OpenStudy (liizzyliizz):

the radius of a sphere is increasing at a constant of 2 cm/sec. at the instant when the volume of the sphere is increasing at 32pi cm^3. sec, the surface area of the sphere is?

OpenStudy (liizzyliizz):

I am not sure where to start in this problem o.o help?

OpenStudy (anonymous):

first we need an equation that relates surface area, radius, and volume

OpenStudy (amistre64):

surface are of sphere = 4pi r right?

OpenStudy (liizzyliizz):

yes it is :)

OpenStudy (amistre64):

or is it 4pi r^2

OpenStudy (liizzyliizz):

r^2

OpenStudy (anonymous):

should be squared since it is an area right?

OpenStudy (anonymous):

i will take your word for the 4

OpenStudy (liizzyliizz):

the second one is correct :P

OpenStudy (amistre64):

so rate of change of SA is: 8 pi r * rate of change of radius

OpenStudy (amistre64):

\[S=4\pi r^2\] \[S' = 8\pir*\frac{dr}{dt}\]

OpenStudy (anonymous):

preview isn't working for me either, for the last few days

OpenStudy (amistre64):

when the rate of change of the volume is: so Volume = 4/3 pi r^3 V' = 4 pi r^2 * dr/dt

OpenStudy (anonymous):

Answer is 16 cm^2

OpenStudy (amistre64):

im thinking it says: when V' = 32 and the dr/dt is 3 that we can find the radius and in turn find the surface area; but surface area IS V' :)

OpenStudy (anonymous):

dr/dt = 2

OpenStudy (anonymous):

yeah i jumped the gun. time for a piece of paper

OpenStudy (anonymous):

dV/dt = S.(dr/dt)

OpenStudy (anonymous):

hence, S = 0.5(dV/dt) = 0.5 *32 = 16

OpenStudy (amistre64):

yeah, trying to parse one long string of information can get challenging at this age: the radius of a sphere is increasing at a constant of 2 cm/sec. at the instant when the volume of the sphere is increasing at 32pi cm^3. sec, the surface area of the sphere is? Instant rate is the derivative; V' which is also the Sa :) 32 = 4 pi r^2 * r' 8/(pi r') = r^2 r = 2 sqrt(2/pi r') r = \(2\sqrt{\frac{2}{2pi}}\) r = \(2\sqrt{\frac{1}{pi}}\) = 2(pi)^(-1/2) that looks fun lol

OpenStudy (anonymous):

god i kept thinking it was 3

OpenStudy (amistre64):

i get 16 too :)

OpenStudy (liizzyliizz):

o.o God help me during the AP exam. lol

OpenStudy (anonymous):

we have \[V'=4\pi r^2 r'\] \[V'=8\pi r^2\] and we know that \[V' = 32 \pi\] so \[8 \pi r^2 = 32\pi\] \[r^24, r=2\]

OpenStudy (anonymous):

that is \[r^2=4\implies r = 2\]

OpenStudy (amistre64):

32pi? ugh

OpenStudy (anonymous):

replace r by 2 in the formula for surface area and get the answer

OpenStudy (anonymous):

yeah pi should not be in answer

OpenStudy (anonymous):

\[S=4\pi \times 2^2 = 16\pi\] what mainaknag said

OpenStudy (anonymous):

problem was not that it was hard, that we both (amistre and myself) got off to a false start because usually the problem says "how fast" not "what is"

OpenStudy (anonymous):

:) I meissed pi in the dV/dt

OpenStudy (amistre64):

i heard the starting pistol and ran the other way :)

OpenStudy (anonymous):

in fact mainaknag solution was snappy two step one

OpenStudy (anonymous):

I think working extra hard for act exams, missing a lot of information in the question

OpenStudy (liizzyliizz):

:O